# Reducing Size Bias in Sampling for Infectious Disease Spread on Networks

### v1

@@ Appendix C Disease metrics
We compare the sampling algorithms (RW and MHRW) based on the accuracy of disease metric estimates for the ER, SW and SF networks using the method described in Section 2.4. Specifically, we discuss the estimates of the time to infection for increasing $\beta$ values while keeping $\gamma=1$ in Section C.1. In Section C.2, we compare the sampling algorithms for three disease metrics: 1) proportion of infected nodes, 2) average number of secondary infections, and 3) time to infection for the ER and SW networks, with $\beta$ values from 00 to $1$ and $\gamma\in\{1/7,1/14\}$.
@@ C.1 Time to get infection
In Figure 8, we compare the estimates of the time to infection from the RW and MHRW samples with the UN for increasing $\beta$ values and $\gamma=1$ for three networks (ER, SW, and SF). In Panel 1 (Figure 7(a), 7(b), 7(c)) results for samples with duplicate nodes and in Panel 2 (Figure 7(d),7(e),7(f)) results for samples without duplicate nodes are shown.
We observe that for ER networks and $\beta<0.2$, there is a large variation in the estimates of both the sampling algorithms in both panels. For $\beta\geq 0.2$, the MHRW sample estimates align well with the UN, while the RW sample estimates slightly underestimate the time to infection, if duplicate nodes are retained in the samples (Figure 7(a)). When duplicate nodes are removed, the MHRW sample estimates slightly underestimate the time to infection (Figure 7(d)), although the behaviour of the RW sample estimates remains consistent. For SW networks, a behaviour similar to that of the ER networks is observed. However, the large variation in the estimates for both sampling algorithms persists up to $\beta=0.5$.
@@
Overall, both sampling algorithms are performing well across network types, particularly for the Scale-free (SF) networks, which is not the case for other two disease metrics (the proportion of infected nodes and the number of secondary infections).
@@ C.2.1 Samples with duplicate nodes
For the ER network, as shown in Figures 8(a) and 8(b), the sample estimates for the proportion of infected nodes become closer to the UN as the recovery rate decreases from $1$ to $1/14$, for both sampling algorithms. Especially, the MHRW sample estimates are well aligned with the UN for $\gamma=1/14$. For the RW sample estimates, there is a significant drop in overestimation with decrease in the recovery rate from $1$ to $1/14$. We observe similar behaviour for the SW networks, as shown in Figures 8(c) and 8(d).
@@
For the estimates of the time to infection in the ER network, as shown in Figures 8(i) and 8(j), the sample estimates deviate away from the UN for both sampling algorithms as recovery rate decreases. The overestimation increases for the MHRW samples and underestimation increases for the RW sample estimates. For the SW networks, as shown in Figures 8(k) and 8(l), underestimation increases for the RW sample estimates. Whereas the MHRW sample estimates align well with the UN with the decrease in recovery rate from $1$ to $1/14$ for $\beta<0.4$.
@@
(3) Time to infection (days).
@@ C.2.2 Samples without duplicate nodes
(3) Time to infection (days).
@@
For the estimates of the time to infection in the ER network, as shown in Figures 9(i) and 9(j), we observe that sample estimates deviate away from the UN for both sampling algorithms as the recovery rate decreases. The underestimation increases for the RW sample estimates and overestimation increases for the MHRW sample estimates. For the SW networks, as shown in Figures 9(k) and 9(l), underestimation increases for the RW sample estimates. Whereas the MHRW sample estimates align well with the UN with the decrease in recovery rate from $1$ to $1/14$ for $\beta<0.4$.

### v2

@@ Appendix C Disease metrics
We compare the sampling algorithms (RW and MHRW) based on the accuracy of disease metric estimates for the ER, SW and SF networks using the method described in Section 2.4. Specifically, we discuss the estimates of the time-to-infection for increasing $\beta$ values while keeping $\gamma=1$ in Section C.1. In Section C.2, we compare the sampling algorithms for three disease metrics: 1) proportion of infected nodes, 2) average number of secondary infections, and 3) time-to-infection for the ER and SW networks, with $\beta$ values from 00 to $1$ and $\gamma\in\{1/7,1/14\}$.
@@ C.1 Time to get infection
In Figure 8, we compare the estimates of the time-to-infection from the RW and MHRW samples with the UN for increasing $\beta$ values and $\gamma=1$ for three networks (ER, SW, and SF). In Panel 1 (Figure 7(a), 7(b), 7(c)) results for samples with duplicate nodes and in Panel 2 (Figure 7(d),7(e),7(f)) results for samples without duplicate nodes are shown.
We observe that for ER networks and $\beta<0.2$, there is a large variation in the estimates of both the sampling algorithms in both panels. For $\beta\geq 0.2$, the MHRW sample estimates align well with the UN, while the RW sample estimates slightly underestimate the time-to-infection, if duplicate nodes are retained in the samples (Figure 7(a)). When duplicate nodes are removed, the MHRW sample estimates slightly underestimate the time-to-infection (Figure 7(d)), although the behaviour of the RW sample estimates remains consistent. For SW networks, a behaviour similar to that of the ER networks is observed. However, the large variation in the estimates for both sampling algorithms persists up to $\beta=0.5$.
@@
Overall, both sampling algorithms are performing well across network types, particularly for the Scale-free (SF) networks, which is not the case for the other two disease metrics (the proportion of infected nodes and the number of secondary infections).
@@ C.2.1 Samples with duplicate nodes
For the ER network, as shown in Figures 8(a) and 8(b), the sample estimates for the proportion of infected nodes become closer to the UN as the recovery rate decreases from $1$ to $1/14$, for both sampling algorithms. Especially, the MHRW sample estimates are well aligned with the UN for $\gamma=1/14$. For the RW sample estimates, there is a significant drop in overestimation with the decrease in the recovery rate from $1$ to $1/14$. We observe similar behaviour for the SW networks, as shown in Figures 8(c) and 8(d).
@@
For the estimates of the time-to-infection in the ER network, as shown in Figures 8(i) and 8(j), the sample estimates deviate away from the UN for both sampling algorithms as recovery rate decreases. The overestimation increases for the MHRW samples and the underestimation increases for the RW sample estimates. For the SW networks, as shown in Figures 8(k) and 8(l), underestimation increases for the RW sample estimates. Whereas the MHRW sample estimates align well with the UN with the decrease in recovery rate from $1$ to $1/14$ for $\beta<0.4$.
@@
(3) time-to-infection (days).
@@ C.2.2 Samples without duplicate nodes
(3) time-to-infection (days).
@@
For the estimates of the time-to-infection in the ER network, as shown in Figures 9(i) and 9(j), we observe that sample estimates deviate away from the UN for both sampling algorithms as the recovery rate decreases. The underestimation increases for the RW sample estimates and overestimation increases for the MHRW sample estimates. For the SW networks, as shown in Figures 9(k) and 9(l), underestimation increases for the RW sample estimates. Whereas the MHRW sample estimates align well with the UN with the decrease in recovery rate from $1$ to $1/14$ for $\beta<0.4$.